# Curated LLM: Synergy of LLMs and Data Curation for tabular augmentation in ultra low-data regimes

### v1

@@ Appendix C Additional results
covid (n=20) 80.47 69.85 76.35 76.73 normal-$\uparrow$ 74.69 75.18 $\uparrow$ 59.99 66.87 $\uparrow$ 68.78 69.84 $\uparrow$ 63.94 71.81 $\uparrow$ 69.28 72.32 $\uparrow$ 68.36 68.44 $\uparrow$ 63.92 70.12 $\uparrow$
@@
maggic (n=20) 70.46 56.55 63.67 64.12 normal-$\uparrow$ 58.74 59.98 $\uparrow$ 53.27 55.37 $\uparrow$ 54.40 55.28 $\uparrow$ 50.39 55.27 $\uparrow$ 56.52 59.00 $\uparrow$ 55.83 56.45 $\uparrow$ 54.59 57.82 $\uparrow$
seer (n=20) 90.91 85.96 88.63 88.74 normal-$\uparrow$ 87.90 88.11 $\uparrow$ 78.22 83.69 $\uparrow$ 83.17 84.74 $\uparrow$ 51.07 77.84 $\uparrow$ 84.86 84.75 85.72 86.11 $\uparrow$ 80.33 82.63 $\uparrow$
@@
adult (n=20) 88.37 79.24 45.28 71.07 $\uparrow$ 50.89 75.77 $\uparrow$ 80.03 81.27 normal-$\uparrow$ 73.19 74.10 $\uparrow$ 68.84 79.66 $\uparrow$ 75.65 80.74 $\uparrow$ 74.61 75.04 $\uparrow$ 76.82 78.84 $\uparrow$
@@
covid (n=40) 80.50 71.20 76.17 76.91 normal-$\uparrow$ 75.32 76.13 $\uparrow$ 64.56 68.82 $\uparrow$ 71.08 72.77 $\uparrow$ 61.40 71.21 $\uparrow$ 70.74 73.84 $\uparrow$ 68.85 69.65 $\uparrow$ 64.13 69.58 $\uparrow$
cutract (n=40) 79.14 71.90 74.05 74.82 normal-$\uparrow$ 72.69 72.89 $\uparrow$ 68.43 70.97 $\uparrow$ 70.67 72.39 $\uparrow$ 61.50 74.10 $\uparrow$ 67.98 71.55 $\uparrow$ 70.26 70.12 67.87 70.88 $\uparrow$
maggic (n=40) 70.36 59.65 61.89 63.39 normal-$\uparrow$ 60.41 61.84 $\uparrow$ 55.81 56.64 $\uparrow$ 56.58 58.08 $\uparrow$ 48.68 58.82 $\uparrow$ 57.50 60.90 $\uparrow$ 57.98 58.00 $\uparrow$ 56.61 59.42 $\uparrow$
seer (n=40) 90.93 87.86 88.50 88.89 normal-$\uparrow$ 85.66 85.62 85.95 87.62 $\uparrow$ 84.92 85.45 $\uparrow$ 60.75 85.84 $\uparrow$ 88.46 88.85 $\uparrow$ 86.56 87.58 $\uparrow$ 86.30 88.36 $\uparrow$
compas (n=40) 72.35 65.50 70.96 70.93 65.70 66.72 $\uparrow$ 60.24 63.78 $\uparrow$ 59.91 62.49 $\uparrow$ 65.17 68.19 $\uparrow$ 62.63 67.00 $\uparrow$ 63.70 64.18 $\uparrow$ 58.63 64.17 $\uparrow$
adult (n=40) 88.65 82.23 44.99 74.25 $\uparrow$ 56.38 76.76 $\uparrow$ 78.24 82.22 $\uparrow$ 71.83 81.15 $\uparrow$ 73.65 81.54 $\uparrow$ 81.49 84.34 normal-$\uparrow$ 81.80 83.16 $\uparrow$ 80.54 83.02 $\uparrow$
drug (n=40) 81.26 71.74 78.98 78.88 71.54 73.21 $\uparrow$ 71.34 74.44 $\uparrow$ 72.12 73.68 $\uparrow$ 70.00 76.45 $\uparrow$ 66.91 73.78 $\uparrow$ 69.09 69.75 $\uparrow$ 68.21 70.94 $\uparrow$
covid (n=100) 80.59 74.19 76.53 77.34 normal-$\uparrow$ 74.74 76.27 $\uparrow$ 72.30 75.15 $\uparrow$ 75.45 76.20 $\uparrow$ 69.05 74.94 $\uparrow$ 71.01 76.38 $\uparrow$ 71.71 72.28 $\uparrow$ 73.90 75.17 $\uparrow$
cutract (n=100) 78.93 76.73 74.62 75.70 $\uparrow$ 73.75 74.40 $\uparrow$ 73.89 75.61 $\uparrow$ 75.70 74.49 60.03 74.35 $\uparrow$ 75.49 76.86 normal-$\uparrow$ 76.13 75.47 73.85 76.21 $\uparrow$
maggic (n=100) 70.48 60.14 61.31 63.64 normal-$\uparrow$ 57.17 60.28 $\uparrow$ 59.32 61.13 $\uparrow$ 58.12 58.97 $\uparrow$ 49.10 59.95 $\uparrow$ 60.43 63.06 $\uparrow$ 59.87 59.59 59.43 61.28 $\uparrow$
seer (n=100) 90.93 88.92 88.16 88.69 $\uparrow$ 88.09 88.25 $\uparrow$ 88.45 89.11 $\uparrow$ 87.53 88.25 $\uparrow$ 83.13 88.44 $\uparrow$ 88.46 89.27 normal-$\uparrow$ 87.95 87.65 86.81 88.67 $\uparrow$
compas (n=100) 72.30 67.75 71.21 71.34 normal-$\uparrow$ 65.61 67.67 $\uparrow$ 64.14 66.72 $\uparrow$ 61.55 65.03 $\uparrow$ 67.43 68.88 $\uparrow$ 65.84 68.39 $\uparrow$ 66.39 66.53 $\uparrow$ 65.61 67.77 $\uparrow$
adult (n=100) 88.42 85.59 46.71 78.68 $\uparrow$ 48.64 75.94 $\uparrow$ 82.30 84.48 $\uparrow$ 77.41 78.98 $\uparrow$ 83.99 84.99 $\uparrow$ 82.67 85.77 normal-$\uparrow$ 85.63 85.43 81.32 83.97 $\uparrow$
drug (n=100) 81.42 74.34 80.35 80.01 70.00 71.76 $\uparrow$ 71.66 75.80 $\uparrow$ 71.26 72.40 $\uparrow$ 74.57 78.65 $\uparrow$ 69.34 77.50 $\uparrow$ 71.04 71.75 $\uparrow$ 71.91 75.38 $\uparrow$
covid (n=200) 80.62 76.73 76.06 77.25 $\uparrow$ 75.18 77.09 $\uparrow$ 76.11 77.54 $\uparrow$ 77.51 78.09 normal-$\uparrow$ 71.92 76.62 $\uparrow$ 74.06 77.53 $\uparrow$ 74.94 74.94 72.75 76.00 $\uparrow$
cutract (n=200) 78.58 77.53 75.28 76.30 $\uparrow$ 74.36 75.38 $\uparrow$ 75.25 75.88 $\uparrow$ 77.66 77.66 75.37 76.69 $\uparrow$ 76.45 77.50 $\uparrow$ 76.77 76.59 72.94 74.86 $\uparrow$
maggic (n=200) 70.28 63.32 61.89 64.30 normal-$\uparrow$ 59.10 62.10 $\uparrow$ 61.57 63.98 $\uparrow$ 56.77 57.35 $\uparrow$ 51.05 63.17 $\uparrow$ 59.17 63.84 $\uparrow$ 62.86 62.86 60.94 63.47 $\uparrow$
seer (n=200) 90.94 90.01 88.65 89.20 $\uparrow$ 87.57 88.14 $\uparrow$ 88.20 88.80 $\uparrow$ 89.69 89.81 $\uparrow$ 89.10 89.68 $\uparrow$ 87.35 89.54 $\uparrow$ 89.13 89.14 $\uparrow$ 88.74 89.65 $\uparrow$
compas (n=200) 72.47 69.14 70.53 71.36 normal-$\uparrow$ 63.49 66.50 $\uparrow$ 66.26 68.73 $\uparrow$ 64.83 66.40 $\uparrow$ 67.85 69.93 $\uparrow$ 60.78 68.07 $\uparrow$ 68.18 68.18 66.26 68.99 $\uparrow$
adult (n=200) 88.50 86.94 38.89 80.35 $\uparrow$ 56.97 79.07 $\uparrow$ 85.25 86.29 $\uparrow$ 86.81 86.63 85.70 85.94 $\uparrow$ 84.18 86.65 $\uparrow$ 86.88 86.70 84.09 85.91 $\uparrow$
drug (n=200) 81.33 77.47 79.06 79.61 $\uparrow$ 71.56 72.54 $\uparrow$ 76.15 78.84 $\uparrow$ 69.00 73.14 $\uparrow$ 77.51 79.72 $\uparrow$ 75.47 80.07 normal-$\uparrow$ 75.82 75.69 71.76 76.95 $\uparrow$
C.2 Full results for performance evaluation
@@
covid (n=40) $75.02_{(0.22)}$ $70.77_{(0.96)}$ $73.40_{(0.61)}$ $\textbf{73.95}_{(0.67)}$ $70.42_{(0.92)}$ $71.93_{(0.60)}$ $63.63_{(1.15)}$ $68.46_{(0.93)}$ $70.50_{(1.49)}$ $70.44_{(1.37)}$ $56.50_{(0.83)}$ $68.68_{(1.39)}$ $66.41_{(2.51)}$ $70.48_{(1.48)}$ $68.66_{(1.48)}$ $68.44_{(1.28)}$ $61.03_{(0.96)}$ $67.35_{(0.60)}$
cutract (n=40) $72.57_{(0.48)}$ $69.18_{(0.65)}$ $69.87_{(0.62)}$ $\textbf{71.72}_{(0.46)}$ $68.47_{(0.59)}$ $69.56_{(0.50)}$ $63.01_{(2.33)}$ $67.87_{(1.36)}$ $65.63_{(2.71)}$ $67.27_{(2.38)}$ $54.39_{(1.59)}$ $68.44_{(0.41)}$ $61.40_{(1.95)}$ $67.98_{(1.06)}$ $67.86_{(0.72)}$ $67.95_{(0.59)}$ $59.79_{(1.85)}$ $66.62_{(1.06)}$

### v3

@@ C.4 Comparison to random noise baseline
covid (n=20) 80.47 69.85 76.35 76.73 $\uparrow$ 74.69 75.18 $\uparrow$ 59.99 66.87 $\uparrow$ 68.78 69.84 $\uparrow$ 63.94 71.81 $\uparrow$ 69.28 72.32 $\uparrow$ 68.36 68.44 $\uparrow$ 63.92 70.12 $\uparrow$
@@
maggic (n=20) 70.46 56.55 63.67 64.12 $\uparrow$ 58.74 59.98 $\uparrow$ 53.27 55.37 $\uparrow$ 54.40 55.28 $\uparrow$ 50.39 55.27 $\uparrow$ 56.52 59.00 $\uparrow$ 55.83 56.45 $\uparrow$ 54.59 57.82 $\uparrow$
seer (n=20) 90.91 85.96 88.63 88.74 $\uparrow$ 87.90 88.11 $\uparrow$ 78.22 83.69 $\uparrow$ 83.17 84.74 $\uparrow$ 51.07 77.84 $\uparrow$ 84.86 84.75 85.72 86.11 $\uparrow$ 80.33 82.63 $\uparrow$
@@
adult (n=20) 88.37 79.24 45.28 71.07 $\uparrow$ 50.89 75.77 $\uparrow$ 80.03 81.27 $\uparrow$ 73.19 74.10 $\uparrow$ 68.84 79.66 $\uparrow$ 75.65 80.74 $\uparrow$ 74.61 75.04 $\uparrow$ 76.82 78.84 $\uparrow$
@@
covid (n=40) 80.47 71.20 76.17 76.91 $\uparrow$ 75.32 76.13 $\uparrow$ 64.56 68.82 $\uparrow$ 71.08 72.77 $\uparrow$ 61.40 71.21 $\uparrow$ 70.74 73.84 $\uparrow$ 68.85 69.65 $\uparrow$ 64.13 69.58 $\uparrow$
cutract (n=40) 79.12 71.90 74.05 74.82 $\uparrow$ 72.69 72.89 $\uparrow$ 68.43 70.97 $\uparrow$ 70.67 72.39 $\uparrow$ 61.50 74.10 $\uparrow$ 67.98 71.55 $\uparrow$ 70.26 70.12 67.87 70.88 $\uparrow$
maggic (n=40) 70.46 59.65 61.89 63.39 $\uparrow$ 60.41 61.84 $\uparrow$ 55.81 56.64 $\uparrow$ 56.58 58.08 $\uparrow$ 48.68 58.82 $\uparrow$ 57.50 60.90 $\uparrow$ 57.98 58.00 $\uparrow$ 56.61 59.42 $\uparrow$
seer (n=40) 90.91 87.86 88.50 88.89 $\uparrow$ 85.66 85.62 85.95 87.62 $\uparrow$ 84.92 85.45 $\uparrow$ 60.75 85.84 $\uparrow$ 88.46 88.85 $\uparrow$ 86.56 87.58 $\uparrow$ 86.30 88.36 $\uparrow$
compas (n=40) 73.02 65.50 70.96 70.93 65.70 66.72 $\uparrow$ 60.24 63.78 $\uparrow$ 59.91 62.49 $\uparrow$ 65.17 68.19 $\uparrow$ 62.63 67.00 $\uparrow$ 63.70 64.18 $\uparrow$ 58.63 64.17 $\uparrow$
adult (n=40) 88.37 82.23 44.99 74.25 $\uparrow$ 56.38 76.76 $\uparrow$ 78.24 82.22 $\uparrow$ 71.83 81.15 $\uparrow$ 73.65 81.54 $\uparrow$ 81.49 84.34 $\uparrow$ 81.80 83.16 $\uparrow$ 80.54 83.02 $\uparrow$
drug (n=40) 81.41 71.74 78.98 78.88 71.54 73.21 $\uparrow$ 71.34 74.44 $\uparrow$ 72.12 73.68 $\uparrow$ 70.00 76.45 $\uparrow$ 66.91 73.78 $\uparrow$ 69.09 69.75 $\uparrow$ 68.21 70.94 $\uparrow$
covid (n=100) 80.47 74.19 76.53 77.34 $\uparrow$ 74.74 76.27 $\uparrow$ 72.30 75.15 $\uparrow$ 75.45 76.20 $\uparrow$ 69.05 74.94 $\uparrow$ 71.01 76.38 $\uparrow$ 71.71 72.28 $\uparrow$ 73.90 75.17 $\uparrow$
cutract (n=100) 79.12 76.73 74.62 75.70 $\uparrow$ 73.75 74.40 $\uparrow$ 73.89 75.61 $\uparrow$ 75.70 74.49 60.03 74.35 $\uparrow$ 75.49 76.86 $\uparrow$ 76.13 75.47 73.85 76.21 $\uparrow$
maggic (n=100) 70.46 60.14 61.31 63.64 $\uparrow$ 57.17 60.28 $\uparrow$ 59.32 61.13 $\uparrow$ 58.12 58.97 $\uparrow$ 49.10 59.95 $\uparrow$ 60.43 63.06 $\uparrow$ 59.87 59.59 59.43 61.28 $\uparrow$
seer (n=100) 90.91 88.92 88.16 88.69 $\uparrow$ 88.09 88.25 $\uparrow$ 88.45 89.11 $\uparrow$ 87.53 88.25 $\uparrow$ 83.13 88.44 $\uparrow$ 88.46 89.27 $\uparrow$ 87.95 87.65 86.81 88.67 $\uparrow$
compas (n=100) 73.02 67.75 71.21 71.34 $\uparrow$ 65.61 67.67 $\uparrow$ 64.14 66.72 $\uparrow$ 61.55 65.03 $\uparrow$ 67.43 68.88 $\uparrow$ 65.84 68.39 $\uparrow$ 66.39 66.53 $\uparrow$ 65.61 67.77 $\uparrow$
adult (n=100) 88.37 85.59 46.71 78.68 $\uparrow$ 48.64 75.94 $\uparrow$ 82.30 84.48 $\uparrow$ 77.41 78.98 $\uparrow$ 83.99 84.99 $\uparrow$ 82.67 85.77 $\uparrow$ 85.63 85.43 81.32 83.97 $\uparrow$
drug (n=100) 81.41 74.34 80.35 80.01 70.00 71.76 $\uparrow$ 71.66 75.80 $\uparrow$ 71.26 72.40 $\uparrow$ 74.57 78.65 $\uparrow$ 69.34 77.50 $\uparrow$ 71.04 71.75 $\uparrow$ 71.91 75.38 $\uparrow$
covid (n=200) 80.47 76.73 76.06 77.25 $\uparrow$ 75.18 77.09 $\uparrow$ 76.11 77.54 $\uparrow$ 77.51 78.09 $\uparrow$ 71.92 76.62 $\uparrow$ 74.06 77.53 $\uparrow$ 74.94 74.94 72.75 76.00 $\uparrow$
cutract (n=200) 79.12 77.53 75.28 76.30 $\uparrow$ 74.36 75.38 $\uparrow$ 75.25 75.88 $\uparrow$ 77.66 77.66 75.37 76.69 $\uparrow$ 76.45 77.50 $\uparrow$ 76.77 76.59 72.94 74.86 $\uparrow$
maggic (n=200) 70.46 63.32 61.89 64.30 $\uparrow$ 59.10 62.10 $\uparrow$ 61.57 63.98 $\uparrow$ 56.77 57.35 $\uparrow$ 51.05 63.17 $\uparrow$ 59.17 63.84 $\uparrow$ 62.86 62.86 60.94 63.47 $\uparrow$
seer (n=200) 90.91 90.01 88.65 89.20 $\uparrow$ 87.57 88.14 $\uparrow$ 88.20 88.80 $\uparrow$ 89.69 89.81 $\uparrow$ 89.10 89.68 $\uparrow$ 87.35 89.54 $\uparrow$ 89.13 89.14 $\uparrow$ 88.74 89.65 $\uparrow$
compas (n=200) 73.02 69.14 70.53 71.36 $\uparrow$ 63.49 66.50 $\uparrow$ 66.26 68.73 $\uparrow$ 64.83 66.40 $\uparrow$ 67.85 69.93 $\uparrow$ 60.78 68.07 $\uparrow$ 68.18 68.18 66.26 68.99 $\uparrow$
adult (n=200) 88.37 86.94 38.89 80.35 $\uparrow$ 56.97 79.07 $\uparrow$ 85.25 86.29 $\uparrow$ 86.81 86.63 85.70 85.94 $\uparrow$ 84.18 86.65 $\uparrow$ 86.88 86.70 84.09 85.91 $\uparrow$
drug (n=200) 81.41 77.47 79.06 79.61 $\uparrow$ 71.56 72.54 $\uparrow$ 76.15 78.84 $\uparrow$ 69.00 73.14 $\uparrow$ 77.51 79.72 $\uparrow$ 75.47 80.07 $\uparrow$ 75.82 75.69 71.76 76.95 $\uparrow$
C.6 Full results for performance evaluation
@@
covid (n=40) $74.41_{(0.11)}$ $70.77_{(0.96)}$ $73.40_{(0.61)}$ $\textbf{73.95}_{(0.67)}$ $70.42_{(0.92)}$ $71.93_{(0.60)}$ $63.63_{(1.15)}$ $68.46_{(0.93)}$ $70.50_{(1.49)}$ $70.44_{(1.37)}$ $56.50_{(0.83)}$ $68.68_{(1.39)}$ $66.41_{(2.51)}$ $70.48_{(1.48)}$ $68.66_{(1.48)}$ $68.44_{(1.28)}$ $61.03_{(0.96)}$ $67.35_{(0.60)}$
cutract (n=40) $72.23_{(0.65)}$ $69.18_{(0.65)}$ $69.87_{(0.62)}$ $\textbf{71.72}_{(0.46)}$ $68.47_{(0.59)}$ $69.56_{(0.50)}$ $63.01_{(2.33)}$ $67.87_{(1.36)}$ $65.63_{(2.71)}$ $67.27_{(2.38)}$ $54.39_{(1.59)}$ $68.44_{(0.41)}$ $61.40_{(1.95)}$ $67.98_{(1.06)}$ $67.86_{(0.72)}$ $67.95_{(0.59)}$ $59.79_{(1.85)}$ $66.62_{(1.06)}$